OpenStudy (anonymous):

If x - 3y = 0 and 5x - y = 14, then 6x - 4y =

OpenStudy (anonymous):

The answer choices are a.-14 b.-7 c.-3 d.1 e.14

OpenStudy (anonymous):

multiply 5x-y=14 with 3 and solve it with x-3y=14 then we get x=3 now substitute it in x-3y=0 then we get y=1 now x=3 and y=1 in 6x-4y 6(3)-4(1)=18-4 =14